# State-insensitive wavelengths for light shifts and photon scattering from Zeeman states

### v1

@@ III.2 Raman scattering
The internal state should not change in the scattering process. Not all states can Raman scatter both polarizations of light, as shown in Fig. 2(c), and so the only possible way to eliminate $m$ dependence from the Raman scattering is for it to be zero. This means we require
@@ III.3 Summary
In summary, three conditions on the dipole matrix elements need to be satisfied for state-independent light-scattering. If they are all met at the same drive frequency for all $m$, then the generalized polarizability tensor is independent of the magnetic state and the field will not be correlated with the atomic state. The conditions are summarized below:
@@
    This condition means that the induced dipole has no component perpendicular to the input field. We thus define the detuning at which this condition is met as $\Delta_{\perp}$.
@@
    Equal dipole amplitude. The sum of the different-handed dipole matrix elements must be independent of $m$
@@
    Meeting this condition means that the parallel component of the induced dipole has the same amplitude for all states. We thus define the detuning at which it is met as $\Delta_{\parallel}$.
@@
IV Detunings to meet state-indepedent scattering conditions
@@ IV.1 Preserved dipole ellipticity
Here, we find the detuning that satisfies Eq. (18). This ensures that the induced dipole has the same polarization as the input field. The sum has three terms, since $F^{\prime}\in\{F-1,F,F+1\}$ for the $D_{2}$ line. The set of three detunings $\{\Delta_{F^{\prime}}\}$ are not independent from each other, and two of them can be defined as differences from the other. We define this detuning from the central $F^{\prime}=F$ line, i.e., $\Delta_{F}\equiv\Delta$, such that $\Delta_{F\pm 1}=\Delta+\zeta_{\pm 1}$, where $\zeta_{\pm 1}=\omega_{F}-\omega_{F\pm 1}$ is the hyperfine splitting. By direct evaluation of the dipole matrix elements, the above sum becomes (see Appendix B for details)
@@
where $J$ and $J^{\prime}$ are the electron angular momentum quantum numbers of the ground and excited states respectively and $I$ is the nuclear angular momentum quantum number. Importantly, Eq. (21) contains no $m$ dependence, and so $\Delta_{\perp}$ is the same for all $m$ states. This occurs because $r^{2}_{m,F,F^{\prime}}-s^{2}_{m,F,F^{\prime}}$ is always linear in $m$, and so it simply factors out. The above equation has two trivial solutions, $\Delta_{\perp}\rightarrow\pm\infty$. These solutions simply describe that if the probe light is infinitely far detuned then all $m$ states have no response, trivially making them respond in the same way. There are also two non-trivial solutions that can be found by rearranging the expression into a quadratic equation in $\Delta_{\perp}$. There are always guaranteed to be two real non-trivial solutions for scattering on the $D_{2}$ line. Details on the form of the solutions and their derivation are provided in Appendix B.
@@ IV.2 Equal dipole amplitude
Here we find the detuning which satisfies Eq. (19). This ensures that the induced dipole parallel to the input field has the same amplitude for all $m$ states. Following the same procedure as above, we expand the sum as three terms with $F^{\prime}\in\{F-1,F,F+1\}$ and then expand the dipole matrix elements. In this case we find that $r^{2}_{m,F,F^{\prime}}+s^{2}_{m,F,F^{\prime}}$ has two contributions: a scalar term and a term that is quadratic in $m$. For the scattering rate to be $m$ independent we thus require that the sum over the terms quadratic in $m$ is zero. As above, since the $m$ dependence is the same for every term, it simply factors out, and any solution that is found works for all $m$ states. The detuning at which the above condition is met for all $m$ states is thus the solution to
@@
This equation again has two trivial solutions, $\Delta_{\parallel}\rightarrow\pm\infty$. To find potential non-trivial solutions, we again rearrange the terms into a quadratic equation. In this case, the coefficient on the $\Delta_{\parallel}^{2}$ term is zero for the $D_{2}$ line, and thus the condition has only one non-trivial solution. Details on the form of the solution and its derivation are provided in Appendix C.
@@ IV.3.2 $\pi$ scattering
Once again, the $m$-dependence takes a familiar form: we require Eq. (23) to be fulfilled. The Raman rate becomes $m$ independent at $\Delta_{\parallel}$. The $m$ independent part is zero at a third detuning condition $\Delta_{\pi}$. The expression for $\Delta_{\pi}$ can be simplified using the zero combination of the three Wigner 6-j symbols we found in the derivation of $\Delta_{\parallel}$ [Eq. (56) in App. C]. For scattering on the $D_{2}$ line it is given as
@@ V Approximating the hyperfine splitting
The hyperfine splitting can be expressed in terms of a sum over different nuclear multipole moments [45, 46]. Typically, the largest contribution is given by the magnetic dipole moment interaction, quantified by the magnetic dipole hyperfine constant $A_{\mathrm{hfs}}$, with the largest corrections arising from the electric quadrupole and magnetic octopole moments, quantified by $B_{\mathrm{hfs}}$ and $C_{\mathrm{hfs}}$ respectively.
@@
The hyperfine splitting in the $D_{2}$ line is thus
@@
Substituting these values into the expressions for $\Delta_{\perp},\Delta_{\parallel}$ and $\Delta_{\pi}$ we find that all three of our detunings have a common root (full details of this derivation are provided in Appendix D). This brings us to the key finding of our paper: for atoms scattering on a $D_{2}$ line, there is a detuning where the generalized polarizability tensor does not correlate the atomic state with the field at all.
@@ VI Optimization of the state-insensitive detuning
We have shown that, for cases where hyperfine structure results only from the nuclear magnetic moment term, the hyperfine structure interval rule guarantees that all three state-independence criteria can be met at a single detuning $\Delta$. However, this interval rule does not strictly hold when higher-order nuclear moments contribute to the hyperfine structure. In this more general situation, and as shown in Table I, we no longer find a single detuning at which all three criteria apply. Nevertheless, to the extent that higher-order nuclear moments contribute only weakly to hyperfine structure, the three state-independent criteria are all approximately fulfilled at the same detuning.
@@
Figure 3: Optimization to find the magic detuning. The overall value of the magic distance defined by Eq. (29), is shown along with the different contributions to that value. Vertical dashed lines show the detunings that meet each condition, while the solid vertical line shows the optimized detuning. The example shown here is a ${}^{87}$b atom in the $F=1$ line driven with linear light.
@@
${}^{6}$i [46] 3/2 2.9 -1.8 0.087 0.58, 2.38 2.41 2.32 2.39 $6.7\times 10^{-4}$
${}^{7}$i [47] 1 6.0 -2.9 0.052 1.45, -15.0 -15.5 -15.0 -15.0 $3.0\times 10^{-5}$
${}^{7}$i [47] 2 9.4 -6.0 0.052 1.55, 9.11 8.92 9.40 9.05 $9.4\times 10^{-4}$
${}^{23}$a [48] 1 -34.3 15.8 0.147 -8.0, 85.1 93.8 85.8 85.5 $1.7\times 10^{-4}$
${}^{23}$a [48] 2 -58.3 34.3 0.147 -9.4, -53.4 -50.3 -58.3 -52.4 $7.9\times 10^{-3}$
${}^{40}$[49] 7/2 33.3 -24.2 0.45 3.5, -73.0 -77.8 -74.0 -73.4 $5.7\times 10^{-4}$
${}^{40}$[49] 9/2 44.1 -33.3 0.45 3.9, 60.6 57.6 64.1 60.1 $3.0\times 10^{-3}$
${}^{85}$b [50] 2 -63.4 29.4 1.03 -7.7, 140.6 227.6 147.9 143.9 $7.2\times 10^{-3}$
${}^{85}$b [50] 3 -120.6 63.4 1.03 -13.4, -118.9 -97.3 -150.8 -113.9 $6.0\times 10^{-2}$
${}^{87}$b [51] 1 -156.9 72.2 0.148 -36.4, 389.4 429.4 392.2 391.2 $1.8\times 10^{-4}$
${}^{87}$b [51] 2 -266.7 156.9 0.148 -42.8, -244.3 -229.9 -266.7 -239.5 $8.0\times 10^{-3}$
${}^{133}$s [52] 3 -201.29 151.22 -0.00981 -25.20, 453.04 452.36 452.90 452.99 $3.0\times 10^{-7}$
${}^{133}$s [52] 4 -251.09 201.29 -0.00981 -25.12, -352.05 -352.50 -351.53 -352.13 $1.9\times 10^{-6}$
Table 1: Properties of alkali atoms, values of detunings that meet certain conditions, and optimized detunings for minimal magic distance. Only one line is presented for ${}^{6}$i as, due to its nuclear spin $I=1$, scattering from the $F=1/2$ level is only via two $F^{\prime}$ levels. Two isotopes of potassium, ${}^{39}$ and ${}^{41}$, are also omitted, as the hyperfine structure of the $D_{2}$ line is close to degenerate and not known to sufficient accuracy to perform these calculations [49, 50].
The dependence of the magic distance with frequency, for ${}^{87}$b, is shown in Fig. 3. Imperfections in the generalized polarizability tensor due to Rayleigh scattering into the perpendicular and parallel mode are minimized at $\Delta_{\perp}$ and $\Delta_{\parallel}$, respectively. Raman scattering of circular light is also minimized at $\Delta_{\parallel}$, while Raman scattering of $\pi$-polarized light is minimized close to $\Delta_{\pi}$. Note that this does not occur exactly at $\Delta_{\pi}$ because that detuning is only where the state-independent fraction of the Raman scattering is zero, as the $m$ dependent part is minimized at $\Delta_{\parallel}$, and so minimizing $\pi$-polarized Raman scattering is a compromise between these two values. This is also the reason why the primary contribution of imperfections to the generalized polarizability tensor is $\pi$-polarized Raman scattering. As such, the overall optimal detuning for this species lies relatively close to $\Delta_{\pi}$. Nevertheless, the magic distance is relatively flat across a fairly broad range of detunings, such that small perturbations from the optimal detuning would not drastically increase the magic distance.
@@ VII Optimized detuning
In this Section, we calculate the relevant detunings for different species of atoms and atomic ions. We focus on alkali atoms and singly-ionized alkaline-earth atoms. In both cases, the single outer electron provides a simple and clean spectrum, and the values of $J$ and $J^{\prime}$ are always such that the above analysis is valid. We optimize for linear polarization, yielding an optimized detuning we define as $\Delta_{\updownarrow}$. It should be noted that the optimized detuning does depend on the input field, but the impact is small. The optimized detuning for circular input field is generally within 1 MHz of $\Delta_{\updownarrow}$, and for elliptical polarizations the optimal detuning varies smoothly between these values.
@@ VII.1 Alkali atoms
Figure 4: The normalized difference between $\Delta_{\perp}$ and $\Delta_{\parallel}$, as a function of the ratio of (electric quadrupole vs magnetic dipole) hyperfine contributions, $|B_{\mathrm{hfs}}/A_{\mathrm{hfs}}|$. $\Delta_{\perp}$ is chosen as the solution closest to $\Delta_{\parallel}$. Generically, when the relative contribution of the electric quadrupole moment is very small then the two detunings are closer to one another. Data is shown for ground state hyperfine manifolds $F=I+1/2$ (+ symbols) and $F=I-1/2$ ($\times$ symbols).
Figure 4: The normalized difference between $\Delta_{\perp}$ and $\Delta_{\parallel}$, as a function of the ratio of (electric quadrupole vs magnetic dipole) hyperfine contributions, $|B_{\mathrm{hfs}}/A_{\mathrm{hfs}}|$. $\Delta_{\perp}$ is chosen as the solution closest to $\Delta_{\parallel}$. Generically, when the relative contribution of the electric quadrupole moment is very small then the two detunings are closer to one another. Data is shown for ground state hyperfine manifolds $F=I+1/2$ (+ symbols) and $F=I-1/2$ ($\times$ symbols).
@@
${}^{43}$${}^{+}$ [53] 3 122.0 -88.1 0.223 14.8, -272.6 -282.3 -274.5 -273.3 $1.5\times 10^{-4}$
${}^{43}$${}^{+}$ [53] 4 159.9 -122.0 0.223 15.8, 216.4 210.3 223.9 215.4 $1.0\times 10^{-3}$
${}^{85}$${}^{+}$ [54] 4 275.0 -315.9 -4.08 36.3, -657.8 -503.2 -605.0 -642.2 $0.023$
${}^{85}$${}^{+}$ [54] 5 154.1 -275.0 -4.08 15.0, 423.5 846.3 213.2 450.2 $0.199$
${}^{87}$${}^{+}$ [54] 4 198.4 -203.0 -2.46 24.0, -461.0 -381.1 -436.5 -453.4 $0.010$
${}^{87}$${}^{+}$ [54] 5 157.0 -198.4 -2.46 14.4, 323.8 439.7 235.5 336.1 0.069
${}^{89}$${}^{+}$ [54] 2 171.9 -77.9 1.07 20.5, -380.2 -643.7 -401.1 -389.7 $8.0\times 10^{-3}$
${}^{89}$${}^{+}$ [54] 3 331.9 -171.9 1.07 36.6, 324.4 263.2 414.9 310.2 0.065
${}^{91}$${}^{+}$ [54] 2 163.3 -115.0 -0.234 28.5, -384.5 -365.4 -381.0 -383.0 $2.1\times 10^{-4}$
${}^{91}$${}^{+}$ [54] 3 200.6 -163.3 -0.234 27.5, 273.2 271.2 275.8 272.8 $7.4\times 10^{-5}$
${}^{135}$${}^{+}$ [55] 1 -167.0 54.0 0.522 -27.9, 403.7 920.2 417.5 413.0 $3.8\times 10^{-3}$
${}^{135}$${}^{+}$ [55] 2 -398.0 167.0 0.522 -56.3, -294.9 -233.2 -398.0 -272.6 0.118
${}^{137}$${}^{+}$ [55] 1 -161.9 34.7 0.727 -18.4, 382.6 -1216.0 404.8 398.4 0.011
${}^{137}$${}^{+}$ [55] 2 -474.1 161.9 0.727 -60.9, -314.9 -221.3 -474.1 -283.9 0.251
${}^{221}$${}^{+}$ [56] 2 681.1 -1136.2 -60.9 231.9, -1946.4 -1073.6 -1589.2 -1833.0 0.084
${}^{221}$${}^{+}$ [56] 3 -1001.8 -681.1 -60.9 -227.8, 623.9 734.9 -1252.3 N/A N/A
${}^{223}$${}^{+}$ [56] 1 751.8 -808.3 15.3 368.7, -2060.2 -1224.5 -1879.5 -1965.4 0.025
${}^{223}$${}^{+}$ [56] 2 -1034.3 -751.8 15.3 -270.0, 720.6 795.2 -1034.3 N/A N/A
@@
Table 2: Properties of singly-ionized alkaline earth ions and values of detunings that meet certain conditions, and optimized detunings for minimal magic distance. All integer hyperfine spin isotopes of alkaline-earth atoms have no nuclear spin, and therefore no hyperfine splitting, while ${}^{133}$${}^{+}$ cannot have a magic detuning due to its nuclear spin of $I=1/2$. For the higher $F$ level for both radium isotopes, the optimization procedure for linear polarization finds no minima between $\Delta_{\perp}$ and $\Delta_{\pi}$, such that it instead finds the trivial $\Delta\rightarrow\infty$ solutions, and so no optimized detuning is obtained.
The detuning conditions share a similar root for all considered species. The most dissimilar are found for ${}^{85}$b, where the $F=2$ line has roots almost 80 MHz apart. The closest are found for ${}^{133}$s, where they are within 1 MHz of each other. As anticipated, this is directly related to the contributions to the hyperfine splitting, as ${}^{85}$b and ${}^{133}$s have the highest and lowest relative electric quadrupole contributions respectively. Figure 4 shows that, generically, the relative detuning difference decreases together with the relative electric quadrupole contribution. The two species that do not perfectly fit this trend are the two fermionic isotopes ${}^{6}$i and ${}^{40}$, where the difference is smaller than the general trend. For these species, the different nuclear spin means the electric quadrupole shift contribution is of a different magnitude even for the same $B_{\mathrm{hfs}}$. Since ${}^{23}$a and ${}^{87}$b have the same ratio of hyperfine contributions and the same nuclear spin, they also have the same detuning difference. For all species, the detuning difference is smaller for the hyperfine level $F=I+1/2$ than for $F=I-1/2$.
The usefulness of these magic wavelengths is likely greatly reduced for lithium experiments, as the detuning is found very close to a resonance. For ${}^{6}$i, the detuning conditions are met less than 1 MHz from resonance with the $F^{\prime}=1/2$ excited state. For ${}^{7}$i, they are met around 6 MHz and 10 MHz from excited state resonances for the $F=1$ and $F=2$ respectively. Therefore, to meet the criterion of essentially no atomic excitation, one would need to use extremely weak laser power. Furthermore, there will be much higher sensitivity to frequency differences between magnetic sublevels.

### v2

@@ III.2 Raman scattering
The Zeeman state should not change in the scattering process. Not all states can Raman scatter both polarizations of light, as shown in Fig. 2(c), and so the only possible way to eliminate $m$ dependence from Raman scattering is for it to be zero. Therefore,
@@ III.3 Summary
In summary, three conditions on the dipole matrix elements need to be satisfied for state-independent light-scattering. If they are all met at the same drive frequency for all $m$, then the generalized polarizability tensor is independent of the Zeeman state and the field will not be correlated with the atomic state. The conditions are summarized below:
@@
    We denote the detuning at which this condition is met $\Delta_{\perp}$, as the induced dipole has no component perpendicular to the input field.
@@
    Equal dipole amplitude. The sum of the different-handed dipole matrix elements must be independent of $m$. This implies
@@
    We denote the detuning at which the condition is met $\Delta_{\parallel}$, as the parallel component of the induced dipole has the same amplitude for all states.
@@
IV Detunings to meet state-independent scattering conditions
@@ IV.1 Preserved dipole ellipticity
Here, we find the detuning that satisfies Eq. (18), which ensures that the induced dipole has the same polarization as the input field. In our considered setup, the sum has three terms with $F^{\prime}\in\{F-1,F,F+1\}$. The set of three detunings $\{\Delta_{F^{\prime}}\}$ are not independent from each other, and two of them can be defined as differences from the other. We define the detuning from the central $F^{\prime}=F$ line, i.e., $\Delta_{F}\equiv\Delta$, such that $\Delta_{F\pm 1}=\Delta+\zeta_{\pm 1}$, where $\zeta_{\pm 1}=\omega_{F}-\omega_{F\pm 1}$ is the hyperfine splitting. By direct evaluation of the dipole matrix elements, the above sum becomes (see Appendix B for details)
@@
where $J$ and $J^{\prime}$ are the electron angular momentum quantum numbers of the ground and excited states, respectively, and $I$ is the nuclear angular momentum quantum number. Importantly, Eq. (21) contains no $m$ dependence, and so $\Delta_{\perp}$ is the same for all $m$ states. This occurs because $r^{2}_{m,F,F^{\prime}}-s^{2}_{m,F,F^{\prime}}$ is always linear in $m$, and so it simply factors out. The above equation has two trivial solutions, $\Delta_{\perp}\rightarrow\pm\infty$. These solutions simply describe that if the probe light is infinitely far detuned then all $m$ states have no response, trivially making them respond in the same way. There are also two non-trivial solutions that can be found by rearranging the expression into a quadratic equation in $\Delta_{\perp}$. For the specific case of scattering on a $D_{2}$ line, they are real. Details on the form of the solutions and their derivation are provided in Appendix B.
@@ IV.2 Equal dipole amplitude
Here we find the detuning that satisfies Eq. (19) and ensures that the induced dipole parallel to the input field has the same amplitude for all $m$ states. Following the same procedure as above, we expand the sum in Eq. (19) as three terms with $F^{\prime}\in\{F-1,F,F+1\}$ and then expand the dipole matrix elements. In this case we find that $r^{2}_{m,F,F^{\prime}}+s^{2}_{m,F,F^{\prime}}$ has two contributions: a scalar term and a term that is quadratic in $m$. For the scattering rate to be $m$ independent we thus require the sum over the terms quadratic in $m$ to be zero. As above, since the $m$ dependence is the same for every term, it simply factors out, and any solution that is found works for all $m$ states. The detuning at which the above condition is met for all $m$ states is thus the solution to
@@
This equation has two trivial solutions, $\Delta_{\parallel}\rightarrow\pm\infty$. To find potential non-trivial solutions, we again rearrange the terms into a quadratic equation. In the case of scattering on a $D_{2}$ line, the coefficient for the $\Delta_{\parallel}^{2}$ term is zero, and thus the condition has only one non-trivial solution. Details on the form of the solution and its derivation are provided in Appendix C.
@@ IV.3.2 $\pi$ scattering
This expression has two parts: one proportional to $m$ and one independent of $m$. The $m$-dependent terms take a familiar form, that of Eq. (23). The Raman rate thus becomes $m$ independent at $\Delta_{\parallel}$. The $m$ independent part is zero at a third detuning condition $\Delta_{\pi}$. For scattering on a $D_{2}$ line, the expression for $\Delta_{\pi}$ can be simplified using the zero combination of the three Wigner 6-j symbols we found in the derivation of $\Delta_{\parallel}$ [Eq. (54) in App. C], and is given as
@@ V Approximating the hyperfine splitting
The hyperfine splitting can be expressed in terms of a sum over different nuclear multipole moments [56, 57]. Typically, the largest contribution is given by the magnetic dipole moment interaction, described by the magnetic dipole hyperfine constant $A_{\mathrm{hfs}}$, with the largest corrections arising from the electric quadrupole and magnetic octopole moments, quantified by $B_{\mathrm{hfs}}$ and $C_{\mathrm{hfs}}$ respectively.
@@
The hyperfine splitting between excited states is thus
@@
Substituting these values into the expressions for $\Delta_{\perp},\Delta_{\parallel}$ and $\Delta_{\pi}$ we find that all three of our detunings have a common root (full details of this derivation are provided in Appendix D). This brings us to the key finding of our paper: for atoms scattering on a $D_{2}$ line, there is a detuning where the generalized polarizability tensor does not correlate the Zeeman state with the field at all.
@@ VI Optimization of the state-insensitive detuning
We have shown that, for cases where hyperfine structure results only from the nuclear magnetic dipole moment, the hyperfine structure interval rule guarantees that all three state-independence criteria can be met at a single detuning. However, this interval rule does not strictly hold when higher-order nuclear moments are included. In this more general situation, and as shown in Table I, we no longer find a single detuning at which all three criteria apply. Nevertheless, to the extent that higher-order nuclear moments contribute only weakly to hyperfine structure, the three state-independent criteria are all approximately fulfilled at the same detuning.
@@
Figure 3: Optimization to find the magic detuning. The overall value of the magic distance defined by Eq. (29), is shown along with the different contributions to that value. Vertical dashed lines show the detunings that meet each condition, while the solid vertical line shows the optimized detuning. The example shown here is a 87Rb atom in the $F=1$ line driven with linear light.
@@
6Li [57] 3/2 2.9 -1.8 0.087 0.58, 2.38 2.41 2.32 2.39 $6.7\times 10^{-4}$
7Li [58] 1 6.0 -2.9 0.052 1.45, -15.0 -15.5 -15.0 -15.0 $3.0\times 10^{-5}$
7Li [58] 2 9.4 -6.0 0.052 1.55, 9.11 8.92 9.40 9.05 $9.4\times 10^{-4}$
23Na [59] 1 -34.3 15.8 0.147 -8.0, 85.1 93.8 85.8 85.5 $1.7\times 10^{-4}$
23Na [59] 2 -58.3 34.3 0.147 -9.4, -53.4 -50.3 -58.3 -52.4 $7.9\times 10^{-3}$
40[60] 7/2 33.3 -24.2 0.45 3.5, -73.0 -77.8 -74.0 -73.4 $5.7\times 10^{-4}$
40[60] 9/2 44.1 -33.3 0.45 3.9, 60.6 57.6 64.1 60.1 $3.0\times 10^{-3}$
85Rb [61] 2 -63.4 29.4 1.03 -7.7, 140.6 227.6 147.9 143.9 $7.2\times 10^{-3}$
85Rb [61] 3 -120.6 63.4 1.03 -13.4, -118.9 -97.3 -150.8 -113.9 $6.0\times 10^{-2}$
87Rb [62] 1 -156.9 72.2 0.148 -36.4, 389.4 429.4 392.2 391.2 $1.8\times 10^{-4}$
87Rb [62] 2 -266.7 156.9 0.148 -42.8, -244.3 -229.9 -266.7 -239.5 $8.0\times 10^{-3}$
133Cs [63] 3 -201.29 151.22 -0.00981 -25.20, 453.04 452.36 452.90 452.99 $3.0\times 10^{-7}$
133Cs [63] 4 -251.09 201.29 -0.00981 -25.12, -352.05 -352.50 -351.53 -352.13 $1.9\times 10^{-6}$
Table 1: Properties of alkali atoms, values of detunings that meet certain conditions, and optimized detunings for minimal magic distance. Only one line is presented for 6Li as, due to its nuclear spin $I=1$, scattering from the $F=1/2$ level is only via two $F^{\prime}$ levels. Two isotopes of potassium, 39K and 41K, are also omitted, as the hyperfine structure of the $D_{2}$ line is close to degenerate and not known to sufficient accuracy to perform these calculations [60, 61].
The dependence of the magic distance with frequency for 87Rb is shown in Fig. 3. Imperfections in the generalized polarizability tensor due to Rayleigh scattering into the perpendicular and parallel mode are minimized at $\Delta_{\perp}$ and $\Delta_{\parallel}$, respectively. Raman scattering of circular light is also minimized at $\Delta_{\parallel}$, while Raman scattering of $\pi$-polarized light is minimized close to $\Delta_{\pi}$. Note that this does not occur exactly at $\Delta_{\pi}$ because that detuning is only where the state-independent fraction of the Raman scattering is zero, as the $m$ dependent part is minimized at $\Delta_{\parallel}$, and so minimizing $\pi$-polarized Raman scattering is a compromise between these two values. This is also the reason why the primary contribution of imperfections to the generalized polarizability tensor is $\pi$-polarized Raman scattering. As such, the overall optimal detuning for this species lies relatively close to $\Delta_{\pi}$. Nevertheless, the magic distance is relatively flat across a fairly broad range of detunings, and small perturbations from the optimal detuning do not drastically increase the magic distance.
@@ VII Optimized detuning
In this Section, we calculate the relevant detunings for different species of atoms and atomic ions. We focus on alkali atoms and singly-ionized alkaline-earth atoms. In both cases, the single outer electron provides a simple and clean spectrum, and the values of $J$ and $J^{\prime}$ are always such that the above analysis is valid. We optimize for linear polarization, yielding an optimized detuning we define as $\Delta_{\updownarrow}$. It should be noted that the optimized detuning does depend on the polarization of the input field, but the impact is small. The optimized detuning for a circular input field is generally within 1 MHz of $\Delta_{\updownarrow}$, and for elliptical polarizations the optimal detuning varies smoothly between these values.
@@ VII.1 Alkali atoms
Figure 4: The normalized difference between $\Delta_{\perp}$ and $\Delta_{\parallel}$, as a function of the ratio of (electric quadrupole vs magnetic dipole) hyperfine contributions, $|B_{\mathrm{hfs}}/A_{\mathrm{hfs}}|$. $\Delta_{\perp}$ is chosen as the solution closest to $\Delta_{\parallel}$. Generically, the two detunings are close to one another if the relative contribution of the electric quadrupole moment is very small. Data is shown for ground state hyperfine manifolds $F=I+1/2$ (+ symbols) and $F=I-1/2$ ($\times$ symbols).
The detuning conditions share a similar root for all considered species. The most dissimilar are found for 85Rb, where the $F=2$ line has roots almost 80 MHz apart. The closest are found for 133Cs, where they are within 1 MHz of each other. As anticipated, this is directly related to the contributions to the hyperfine splitting, as 85Rb and 133Cs have the highest and lowest relative electric quadrupole contributions, respectively. Figure 4 shows that, generically, the relative detuning difference decreases together with the relative electric quadrupole contribution. The two species that do not perfectly fit this trend are the two fermionic isotopes 6Li and 40K, where the difference is smaller than the general trend. For these species, the different nuclear spin means the electric quadrupole shift contribution is of a different magnitude even for the same $B_{\mathrm{hfs}}$. Since 23Na and 87Rb have the same ratio of hyperfine contributions and the same nuclear spin, they also have the same detuning difference. For all species, the detuning difference is smaller for the hyperfine level $F=I+1/2$ than for $F=I-1/2$.
The usefulness of these magic wavelengths is likely greatly reduced for lithium experiments, as the detuning is found very close to a resonance. For 6Li, the detuning conditions are met less than 1 MHz from resonance with the $F^{\prime}=1/2$ excited state. For 7Li, they are met around 6 MHz and 10 MHz from excited state resonances for the $F=1$ and $F=2$, respectively. Therefore, to meet the criterion of essentially no atomic excitation, one would need to use extremely weak laser power. Furthermore, there will be much higher sensitivity to frequency differences between Zeeman states.
@@
43Ca+ [64] 3 122.0 -88.1 0.223 14.8, -272.6 -282.3 -274.5 -273.3 $1.5\times 10^{-4}$
43Ca+ [64] 4 159.9 -122.0 0.223 15.8, 216.4 210.3 223.9 215.4 $1.0\times 10^{-3}$
85Sr+ [65] 4 275.0 -315.9 -4.08 36.3, -657.8 -503.2 -605.0 -642.2 $0.023$
85Sr+ [65] 5 154.1 -275.0 -4.08 15.0, 423.5 846.3 213.2 450.2 $0.199$
87Sr+ [65] 4 198.4 -203.0 -2.46 24.0, -461.0 -381.1 -436.5 -453.4 $0.010$
87Sr+ [65] 5 157.0 -198.4 -2.46 14.4, 323.8 439.7 235.5 336.1 0.069
89Sr+ [65] 2 171.9 -77.9 1.07 20.5, -380.2 -643.7 -401.1 -389.7 $8.0\times 10^{-3}$
89Sr+ [65] 3 331.9 -171.9 1.07 36.6, 324.4 263.2 414.9 310.2 0.065
91Sr+ [65] 2 163.3 -115.0 -0.234 28.5, -384.5 -365.4 -381.0 -383.0 $2.1\times 10^{-4}$
91Sr+ [65] 3 200.6 -163.3 -0.234 27.5, 273.2 271.2 275.8 272.8 $7.4\times 10^{-5}$
135Ba+ [66] 1 -167.0 54.0 0.522 -27.9, 403.7 920.2 417.5 413.0 $3.8\times 10^{-3}$
135Ba+ [66] 2 -398.0 167.0 0.522 -56.3, -294.9 -233.2 -398.0 -272.6 0.118
137Ba+ [66] 1 -161.9 34.7 0.727 -18.4, 382.6 -1216.0 404.8 398.4 0.011
137Ba+ [66] 2 -474.1 161.9 0.727 -60.9, -314.9 -221.3 -474.1 -283.9 0.251
221Ra+ [67] 2 681.1 -1136.2 -60.9 231.9, -1946.4 -1073.6 -1589.2 -1833.0 0.084
221Ra+ [67] 3 -1001.8 -681.1 -60.9 -227.8, 623.9 734.9 -1252.3 N/A N/A
223Ra+ [67] 1 751.8 -808.3 15.3 368.7, -2060.2 -1224.5 -1879.5 -1965.4 0.025
223Ra+ [67] 2 -1034.3 -751.8 15.3 -270.0, 720.6 795.2 -1034.3 N/A N/A
@@
Table 2: Properties of singly-ionized alkaline earth ions and values of detunings that meet certain conditions, and optimized detunings for minimal magic distance. All integer hyperfine spin isotopes of alkaline-earth atoms have no nuclear spin, and therefore no hyperfine splitting, while 133Ba+ cannot have a magic detuning due to its nuclear spin of $I=1/2$. For the higher $F$ level for both radium isotopes, the optimization procedure for linear polarization finds no minima between $\Delta_{\perp}$ and $\Delta_{\pi}$, such that it instead finds the trivial $\Delta\rightarrow\infty$ solutions, and so no optimized detuning is obtained.